OpenStudy (gabylovesyou):

The table below shows the ticket rates for whale watching trips offered by Orca Magic Tours.

OpenStudy (gabylovesyou):

Age Ticket Price Under three years free Three to twelve years $32 Over twelve years $44

OpenStudy (gabylovesyou):

On a certain day, the company took 82 people on whale watching trips. There were 54 children aged twelve and under, of which some children were under three years. If x represents the number of children under three years, which equation can be used to find the value of x, where C represents the total amount of money collected from tickets that day?

OpenStudy (gabylovesyou):

(54 + x)32 + (82 – (54 + x)) ∙ 44 = C (54 – x)32 + (82 – 54) ∙ 44 = C (82 + x)32 + (82 – (54 + x)) ∙ 44 = C (82 - x)32 + (82 – 54) ∙ 44 = C

OpenStudy (anonymous):

0x + (54 - x )32 + (82-54)44 = C

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

simplified form, 2960 - 32x = C

OpenStudy (gabylovesyou):

its the second one??

OpenStudy (gabylovesyou):

@sauravshakya

OpenStudy (anonymous):

Yeah @Gabylovesyou

OpenStudy (gabylovesyou):

k thank u!! :)

OpenStudy (anonymous):

WELCOME